If leah is 6 years older than sue, and john is 5 years older than leah, and the total of their ages is 41. Then how old is sue?
Answer by jorel1380(3719)

John=Leah+5
Leah=Sue+6
Leah+Sue+John=41
-----------------------
John=Sue+11
Sue+6+Sue+Sue+11=41
3Sue=24
Sue=8.
Sue is 8 years old.